OpenStudy (anonymous):

identify the tables below on whether the relationship is linear,quadratic,exponential,or inverse variation x y 1 2 2 4 3 8 4 16

OpenStudy (141823):

We know this is an exponential equation because of the increase in the y values. 2 * 2 = 4 4 * 2 = 8 8 * 2 = 16 and so on. Need anything else?

OpenStudy (anonymous):

thank you @141823 =) if you dont mind could you help me identify the others? this just happens to be something i dont understand at all.

OpenStudy (141823):

Yes, I can explain everything except inverse variation. I haven't seen that yet. What grade math are you taking right now btw?

OpenStudy (anonymous):

okay thank you. Im in 9th grade algebra

OpenStudy (141823):

Alright. Got the questions? And maybe after you can teach me a little about those inverse variations lol

OpenStudy (anonymous):

x y 1 -2 2 4 3 10 4 16 x y 1 12 2 6 3 4 4 3 x y 1 16 2 13 3 8 4 1

OpenStudy (141823):

Ok, for the first one we can see the y values increase by 6 -2 + 6 = 4 4 + 6 = 10 10 + 6 = 16 When the first values increase at a constant rate this is linear. Not to be confused with quadratics, that have the second difference increase at a constant rate. Sounds a little familiar?

OpenStudy (anonymous):

okay that makes sense now. i had first put it down as quadratic